# Hear What Matters! Text-conditioned Selective Video-to-Audio Generation

### v1

@@ Abstract
This work introduces a new task, text-conditioned selective video-to-audio (V2A) generation, which produces only the user-intended sound from a multi-object video. This capability is especially crucial in multimedia production, where audio tracks are handled individually for each sound source for precise editing, mixing, and creative control. However, current approaches generate single source-mixed sounds at once, largely because visual features are entangled, and region cues or prompts often fail to specify the source. We propose SelVA, a novel text-conditioned V2A model that treats the text prompt as an explicit selector of target source and modulates video encoder to distinctly extract prompt-relevant video features. The proposed supplementary tokens promote cross-attention by suppressing text-irrelevant activations with efficient parameter tuning, yielding robust semantic and temporal grounding. SelVA further employs a self-augmentation scheme to overcome the lack of mono audio track supervision. We evaluate SelVA on VGG-MonoAudio, a curated benchmark of clean single-source videos for such a task. Extensive experiments and ablations consistently verify its effectiveness across audio quality, semantic alignment, and temporal synchronization. Code and demo are available at https://jnwnlee.github.io/selva-demo/. Work done during an internship at NAVER AI Lab.
@@ 1 Introduction
In a bustling café, you can effortlessly tune into a friend’s laughter amid the chatter, or pick out the sound of a violin from an entire orchestra. This effortless segregation of sounds, achieved through auditory scene analysis, is a hallmark of human perception [audio_scene_analysis]. At the core of this process lies selective attention, which enables us to focus on a specific sound source while filtering out irrelevant noise. Such an attention-driven mechanism allows humans to extract what truly matters from a rich and noisy world.
@@
Recent advances in neural audio generation have enabled the synthesis of realistic sounds from text descriptions or visual scenes in film and game post-production, known as Foley [oh2023demand]. Video-to-audio (V2A) models [rewas, diff-foley, video-foley, mmaudio, multifoley] now generate temporally coherent, context-aware audio directly from visual content. Since they produce a single holistic soundtrack at a time, it is still challenging to produce a user-intended source only. This limitation stands in sharp contrast to the sound production [oh2023demand, challengetta], where sound designers do not sonify every visible object. Instead, they build scenes by layering individually crafted tracks, selecting which elements to include, and then mix and master them, to achieve precise creative control. However, current generative systems often limit such interactivity; even minor omissions in the output require re-synthesizing the entire audio, hindering practical usability.
@@
Figure 1: SelVA turns text prompts into precise selectors of sound sources within a video. The text-conditioned video encoder extracts intent-focused video features that condition the generator to synthesize only the user-specified sound source (e.g., ‘cat meowing’ vs. ‘dog barking’).
@@
In this work, we tackle the selective sound generation problem: generating only the target sound that aligns with a user’s intention, given multimodal cues such as video context and text description. However, this problem poses fundamental challenges. First, grounding of multiple co-occurring sound sources in unconstrained auditory scenes is fundamentally challenging, especially in the absence of explicit source-level supervision. A few works [saganet, hearyourclick] have explored exploiting off-the-shelf segmentation models for visual scenes to localize potential sound sources, under the assumption that visually separable objects emit independent sounds. Yet, these approaches inherently operate on isolated spatial regions and thus fail to capture the broader visual context necessary for sound generation. As a result, they struggle to handle non-object sound sources, such as environmental or diffuse sounds (e.g., rainfall or wind), which cannot be localized to discrete visual boundaries. In addition, incorporating a large segmentation network increases computational overhead and significantly slows down inference.
@@
We propose a novel SelVA method for text-conditioned selective video-to-audio generation. To this end, our SelVA consists of two main modules: (1) a text-conditioned video encoder, and (2) a multimodal-conditioned generator using both text and visual cues. Unlike previous works that learn only generative models with frozen visual-specific encoders, SelVA efficiently modulates the video encoder using text prompts as explicit selectors of audible semantics, implemented through a lightweight cross-attention block specialized with learnable token [SUP]. To achieve a precise cross-modal grounding, the proposed [SUP] token, inspired by selective attention mechanisms in human perception, stabilizes visual-text alignment by mitigating high-norm artifacts  [vit_register, cavmae_sync] in which models tend to highlight irrelevant tokens in attention blocks due to the spurious correlation. Interestingly, [SUP] suppresses irrelevant token activations and strengthens text-conditioned video representations. The training process further employs a self-augmentation strategy that automatically mixes video scenes and conditions the model with targeted text prompts, without explicitly source-separated groundtruth audios.
@@
To assess performance on such a novel task, we introduce VGG-MonoAudio, a new evaluation benchmark comprising videos with a clear visual sound source corresponding to a single-source audio track. Experimental results demonstrate that SelVA achieves state-of-the-art (SoTA) performance on VGG-MonoAudio, showing robustness in terms of audio quality, semantic alignment, and temporal alignments. Our contributions are summarized:
  • SelVA introduces the first framework for text-conditioned selective video-to-audio generation, efficiently learning a text-guided visual encoder and a multimodal-conditioned sound generator to achieve implicit sound-source feature grounding.
@@
  • A learnable supplementary token mitigates spurious cross-modal correlations by refining attention toward intent-relevant cues, and a self-augmentation scheme enables selective learning without costly supervision.
@@
  • Comprehensive experiments with SelVA evaluated on the new VGG-MonoAudio benchmark demonstrate superior audio quality, semantic fidelity, and temporal synchronization performance over existing SoTAs.
@@ 2 Related Work
Cross-modal neural audio generation has been widely explored due to its applicability in multimedia content production. Text-to-audio (T2A) aims to generate audio from an input text prompt, which usually describes the global semantics such as sound sources and their nuanced timbre (e.g., ‘drill buzzing harshly’). The common baseline is first to extract a text embedding from a pretrained text encoder such as CLAP [clap_ms, clap_laion] and T5 [flant5], then use it as a condition of generative models, including diffusion [audioldm, make-an-audio, stableaudio], auto-regressive modeling [audiogen], and flow-matching [tangoflux]. While text prompts offer intuitive semantic control, they inherently lack the ability to convey temporal dynamics of intensity or harmonics in audio [t-foley, fusionnet, sketch2sound]. Meanwhile, video-to-audio (V2A) [mmaudio, moviegen, multifoley, diff-foley, condfoleygen, specvqgan] resolves this issue by generating audio in synchrony with video. Such synchronization entails two complementary goals: semantic and temporal alignment. As a spatiotemporal modality, video conveys rich cues about sounding objects, including appearance, spatial location, and dynamic motion. In practice, current V2A frameworks remain strongly dependent on pretrained visual encoders [clip, tsn, vivit, diff-foley, synchformer] to provide the conditioning representations.
Recently, some works [sonicvisionlm, video-foley, rewas] have leveraged the capability of pretrained T2A models for V2A generation to reduce the training cost and ensure controllability. Most approaches [vatt, mmaudio, moviegen, audiogen-omni] naively hypothesize that complementary relations of video and text conditions, producing high-fidelity audio. Text prompts complement video embeddings by supplying semantic cues that visual encoders often miss (e.g., visual ambiguity such as occlusion of sounding objects caused by camera work) [v2a-occlusion, vatt]. For example, ReWaS [rewas] and Video-Foley [video-foley] rely on text to control the semantics of sound, while Multifoley [multifoley] leveraged text to change the sound timbre. VinTAGe [vintage] generates both on-screen sound from visual cues and off-screen sound from textual cues. However, existing works do not use text to specify which sound sources should be heard. Instead, text serves merely as an auxiliary cue, not to modulate the given visual information. In this paper, we unlock the potential of text prompts by repositioning them as a direct modulator of video embeddings for controllable V2A.
@@
Selective sound generation has only recently begun to emerge for professional multimedia production, where models synthesize audio exclusively for target sound sources. Hayakawa et al[negativeaudioguidance] proposed an iterative, track-wise approach with sequential generation, where sounds produced in previous steps are excluded from the current one. This is accomplished using negative audio guidance that steers the flow-matching process to avoid regenerating audio from prior stages. While the motivation is related to ours, their method heavily relies on the limited separation capability of the pretrained V2A model, especially at the first generation stage. Otherwise, some works [saganet, hearyourclick] have utilized visual region-level cues, such as segmentation masks produced by pretrained models (e.g., SAM2 [sam2]), for object-focused sound generation. However, those approaches have notable limitations in that they necessitate the integration of computationally expensive pretrained segmentation models [sam2, deva], which have often struggled with occluded objects or incorrectly identifying non-object sounding sources (e.g., rain drop, wind blowing). To overcome those limitations, our SelVA is the first to introduce a text prompt for describing the target sound source within the input video for robust selective sound generation.
@@ 3.1 Motivation and problem statement
While existing works [mmaudio, diff-foley, rewas] have achieved promising results in generating holistic sounds aligned with the input video, they often suffer from low fidelity to the text prompts. Specifically, the model occasionally produces undesired outputs, such as off-screen sounds absent from the video, or non-target objects’ audio that appear in the video but are not specified by the text prompt. This limitation arises mainly because most approaches directly feed video features extracted from a frozen visual encoder—typically pretrained for general recognition tasks—into the generation pipeline. Such visual features tend to be noisy and entangled, containing both irrelevant visual cues alongside sound-related semantics. As a result, selectively generating only the intended sound remains challenging. This hinders users from creating harmonious audio in real-world scenarios [oh2023demand, challengetta, negativeaudioguidance]. For example, a professional audio creator often needs to synthesize a soundtrack with various elements such as speech, music, and sound effects under separate controllable conditions. At this juncture, we argue that the text-conditioned video feature grounding could make a huge room for the controllability of V2A.
@@ 3.2 Text-guided visual feature generation
SelVA modulates visual features to encode sound-source relevant information that the text prompt describes. Most V2A models [specvqgan, condfoleygen, foleycrafter, v-aura, mmaudio, multifoley, vaflow] rely on a pretrained vision encoder and, optionally, a text encoder to extract conditioning features. The vision encoders are generally frozen during the training process, serving as visual feature extractors for audio generation. The extracted visual features encompass the global scene context, yet they inherently carry noisy and excessive irrelevant information. Thereby, it impedes the generation of user-intended sound.
@@
We expect hidden video embeddings to contain semantics that are exclusively aligned with text. However, a simple cross-attention mechanism yields a suboptimal result where the model still suffers from generating sounds corresponding to the motion dynamics of non-target instances. For example, when a dog is barking beside a cat, it incorrectly produces a meowing sound, reflecting semantic confusion between co-occurring sound sources. To mitigate this issue, we introduce a novel technique to attach learnable supplementary tokens  [SUP] preceding text embeddings. As shown in Fig. 3 (‘w/o [SUP]’), such artifacts emerge as semantic patch tokens (particularly those associated with motion dynamics) become high-norm outliers in the attention space. However, as in prior approaches on vision transformers [vit_register, cavmae_sync], adding extra tokens to the sequence of visual embeddings increases the computational cost across all encoder blocks. Furthermore, we should design to learn intent-focused visual representations that emphasize regions and cues relevant to the user’s specified sound source. To this end,  [SUP] are simply prepended to text features as follows:
@@ 3.3 Selective sound generation
Our generator, $\mathcal{G}$, adopts a multimodal diffusion transformer (MM-DiT) architecture [sd3, flux]:
@@ 3.4 Training
Auto-mixing video
As input videos usually comprise multiple sound sources, without explicit annotations for each sound source, it is nontrivial to isolate visual features. To address this, we introduce a self-supervised strategy based on video self-augmentation, motivated by audiovisual separation works [ephrat2018looking, lee2021looking], but reformulated for selective V2A generation. Mixed videos are conditioned on text prompts rather than audio, and the model (i.e., video encoder) is optimized for a continuous latent regression instead of predicting a discrete separation mask in prior works. Concretely, two videos are randomly selected, and horizontally concatenated with a random ratio to make the desired {mixed-video, audio, text} pairs. One of the audio-text pairs is randomly chosen to serve as the target. Formally, an input video $V$ in a mini-batch consists of randomly selected two videos $\{V_{\text{tar}},V_{\text{pair}}\}\in\mathbb{R}^{H\times W}$:
@@
Learning conditional feature extraction from mixed sources and multiple conditioned audio generation simultaneously is inherently complex, as both modules depend on each other’s evolving representations [he2020momentum]. To ensure efficient and stable optimization, the joint training is organized into a two-stage learning scheme, allowing each module to converge toward a consistent representation before mutual conditioning. In the first stage, the video encoder learns to understand text prompts from output features of the teacher model [zhang2019your, noisy_student]. In the left part of Fig. 2, while the teacher model $\mathcal{F}_{T}$ (i.e., pretrained Synchformer [synchformer]) takes a single source video $V_{\text{tar}}$ to generate a pseudo feature $\mathbf{v}_{\text{tar}}$, the student model $\mathcal{F}_{S}$ produces a text-guided visual feature from the mixed source. Formally, the video feature $\mathbf{v}$ of student encoder $\mathcal{F}_{S}$ is learned to minimize the L2-norm regression loss:
@@
Next, we train a generator $\mathcal{G}$ while keeping the video encoder frozen in the second stage. We start from the MM-DiT in MMAudio [mmaudio] as the baseline of the generator. Rather than finetuning the whole parameters, we focus on specific modules that handle video features explicitly, as illustrated in the right part of Fig. 2. Specifically, we finetune two sub-modules only: (1) the initial projection layer of the video feature branch (i.e., $W_{\mathbf{v}}$) and (2) the adaptive LayerNorm (adaLN) module of the audio latent branch of multimodal and single-modal transformer blocks (i.e., $W_{\gamma},W_{\beta}$). The model is trained with conditional flow-matching (CFM) [cfm, cfm_ot]. Given a noise distribution $q(\mathbf{a}_{0})\sim\mathcal{N}(\mathbf{0},I)$, training data distribution $q(\mathbf{a}_{1}=\mathbf{a}_{\text{tar}},\mathbf{c})$ with input condition features $\mathbf{c}=(\hat{\mathbf{v}}_{\text{tar}}=\mathcal{F}_{S}(V,\mathbf{t}_{\text{tar}}),\mathbf{t}_{\text{tar}})$, and timestep $t\in[0,1]$, the CFM objective is formulated as:

### v2

@@ Abstract
This work introduces a new task, text-conditioned selective video-to-audio (V2A) generation, which produces only the user-intended sound from a multi-object video. This capability is especially crucial in multimedia production, where audio tracks are handled individually for each sound source for precise editing, mixing, and creative control. We propose SelVA, a novel text-conditioned V2A model that treats the text prompt as an explicit selector to distinctly extract prompt-relevant sound-source visual features from the video encoder. To suppress text-irrelevant activations with efficient video encoder finetuning, the proposed supplementary tokens promote cross-attention to yield robust semantic and temporal grounding. SelVA further employs an autonomous video-mixing scheme in a self-supervised manner to overcome the lack of mono audio track supervision. We evaluate SelVA on VGG-MonoAudio, a curated benchmark of clean single-source videos for such a task. Extensive experiments and ablations consistently verify its effectiveness across audio quality, semantic alignment, and temporal synchronization. Work partly done during an internship at NAVER AI Lab.
@@ 1 Introduction
In a bustling café, you can effortlessly tune into a friend’s laughter amid the chatter, or pick out the sound of a violin from an entire orchestra. This effortless segregation of sounds, achieved through auditory scene analysis, is a hallmark of human perception [1]. At the core of this process lies selective attention, which enables us to focus on a specific sound source while filtering out irrelevant noise. Such an attention-driven mechanism allows humans to extract what truly matters from a rich and noisy world.
@@
Recent advances in neural audio generation have enabled realistic sound synthesis from text descriptions or visual scenes in film and game post-production, known as Foley [50]. Video-to-audio (V2A) models [32, 49, 42, 8, 7] now generate temporally coherent, context-aware audio directly from visual content. However, they typically produce a single holistic soundtrack at a time. This limitation stands in sharp contrast to the sound production [50, 43], where sound designers do not sonify every visible object. Instead, they build scenes by layering individually crafted tracks, selectively including or suppressing sound elements, and then mixing and mastering them, to achieve precise control. However, with current V2A models, even minor omissions in the output require re-synthesizing the entire audio, hindering practical usability.
@@
Figure 1: SelVA turns text prompts into precise selectors of sound sources within a video. The intent-focused video feature conditions the generator to synthesize only the user-specified sound source (e.g., ‘cat meowing’ vs. ‘dog barking’).
@@
In this work, we tackle the selective sound generation problem: generating only the target sound that aligns with a user’s intention, given multimodal cues such as video and text. However, the presence of multiple co-occurring sound sources in real-world videos makes it inherently difficult to isolate the target sound without explicit supervision. A few works [60, 45] have attempted to localize sound sources using off-the-shelf segmentation models, where visual prompts such as points or masks are used to specify the target object. Yet, such approaches inherently operate on isolated spatial regions. Therefore, they struggle to handle broader visual context, including environmental or diffuse sounds (e.g., rainfall or wind), which cannot be localized to discrete visual boundaries. In addition, reliance on large segmentation networks increases substantial computational overhead. In contrast, we propose a new formulation of text-conditioned selective video-to-audio generation, where the target sound source is specified purely through a text prompt. This removes the dependency on explicit spatial prompts while enabling more flexible and semantically expressive control over sound generation.
@@
We propose a novel SelVA method, consisting of two main modules: (1) a text-conditioned video encoder, and (2) a multimodal-conditioned selective audio generator. Unlike previous works [8, 32] that freeze visual-specific encoders, SelVA efficiently trains the video encoder to use text prompts as explicit selectors of audible semantics. To achieve a precise cross-modal grounding, we implement a cross-attention block in the video encoder with the proposed learnable token [SUP]. Inspired by the selective attention mechanism in human perception, [SUP] mitigates high-norm artifacts  [14, 2] in which models tend to highlight irrelevant tokens in attention blocks due to the spurious correlation. To train SelVA without explicitly source-separated groundtruth audios, we further employ an autonomous video-mixing strategy in a self-supervised manner, where two videos are spatially concatenated as input and the audio from one of them is used as the target.
@@
To assess performance on such a novel task, we introduce VGG-MonoAudio, a new evaluation benchmark comprising videos with a clear visual sound source corresponding to a mono audio track. Experimental results demonstrate that SelVA achieves state-of-the-art (SoTA) performance on VGG-MonoAudio, showing robustness in terms of audio quality, semantic alignment, and temporal alignment. Our contributions are summarized:
  • SelVA is the first text-conditioned selective video-to-audio framework that relies solely on text prompts, without spatial guidance.
@@
  • A learnable supplementary token mitigates spurious cross-modal correlations while empowering intent-relevant cues, and the proposed video-mixing scheme enables selective learning without costly supervision.
@@
  • Experiments on new VGG-MonoAudio demonstrate the effectiveness of SelVA, showing superior audio quality, semantic fidelity, and temporal synchronization performance over existing SoTAs.
@@ 2 Related Work
Cross-modal neural audio generation has been widely explored due to its applicability in multimedia content production. Text-to-audio (T2A) aims to generate audio from an input text prompt, which usually describes the global semantics such as sound sources and their nuanced timbre (e.g., ‘drill buzzing harshly’). The common baseline is first to extract a text embedding from a pretrained text encoder such as CLAP [16, 65] and T5 [10], then use it as a condition of generative models, including diffusion [47, 28, 19], auto-regressive modeling [38], and flow-matching [29]. While text prompts offer intuitive semantic control, they inherently lack the ability to convey temporal dynamics of intensity or harmonics in audio [12, 24, 20]. Meanwhile, video-to-audio (V2A) [8, 53, 7, 49, 15, 30] resolves this issue by generating audio in synchrony with video. Such synchronization entails two complementary goals: semantic and temporal alignment. As a spatiotemporal modality, video conveys rich cues about sounding objects, including appearance, spatial location, and dynamic motion. In practice, current V2A frameworks remain strongly dependent on pretrained visual encoders [54, 63, 3, 49, 31] to provide the conditioning representations.
Recently, some works [67, 42, 32] have leveraged the capability of pretrained T2A models for V2A generation to reduce the training cost and ensure controllability. Most approaches [48, 8, 53, 62] naively hypothesize that complementary relations of video and text conditions, producing high-fidelity audio. Text prompts complement video embeddings by supplying semantic cues that visual encoders often miss (e.g., visual ambiguity such as occlusion of sounding objects caused by camera work) [27, 48]. For example, ReWaS [32] and Video-Foley [42] rely on text to control the semantics of sound, while Multifoley [7] leveraged text to change the sound timbre. VinTAGe [39] generates both on-screen sound from visual cues and off-screen sound from textual cues. However, existing works do not use text to specify which sound sources should be heard. Instead, text serves merely as an auxiliary cue, not to modulate the given visual information. In this paper, we unlock the potential of text prompts by repositioning them as a direct modulator of video embeddings for controllable V2A.
@@
Selective sound generation has only recently begun to emerge for professional multimedia production, where models synthesize audio exclusively for target sound sources. Hayakawa et al. [25] proposed an iterative, track-wise approach with sequential generation, where sounds produced in previous steps are excluded from the current one. This is accomplished using negative audio guidance that steers the flow-matching process to avoid regenerating audio from prior stages. While the motivation is related to ours, their method heavily relies on the limited separation capability of the pretrained V2A model, especially at the first generation stage. Otherwise, some works [60, 45] have utilized visual region-level cues, such as segmentation masks produced by pretrained models (e.g., SAM2 [55]), for object-focused sound generation. However, those approaches have notable limitations in that they necessitate the integration of computationally expensive pretrained segmentation models [55, 9], which have often struggled with occluded objects or incorrectly identifying non-object sounding sources (e.g., rain drop, wind blowing). To overcome those limitations, our SelVA is the first to introduce a text prompt for describing the target sound source within the input video for robust selective sound generation.
@@ 3.1 Motivation and problem statement
While existing works [8, 49, 32] have achieved promising results in generating holistic sounds aligned with the input video, they often suffer from low fidelity to the text prompts. Specifically, the model occasionally produces undesired outputs, i.e., non-target objects’ audio that appear in the video but are not specified by the text prompt. This limitation arises mainly because most approaches directly feed video features extracted from a frozen visual encoder—typically pretrained for general recognition tasks—into the generation pipeline. Such visual features tend to be noisy and entangled, containing both irrelevant visual cues alongside sound-related semantics. As a result, selectively generating only the intended sound remains challenging. This hinders users from creating harmonious audio in real-world scenarios [50, 43, 25]. For example, a professional audio creator often needs to synthesize a soundtrack with various elements such as speech, music, and sound effects under separate controllable conditions. At this juncture, we argue that the text-conditioned video feature grounding could make a huge room for the controllability of V2A.
@@ 3.2 Text-guided visual feature generation
SelVA modulates visual features to encode sound-source relevant information that the text prompt describes. Most V2A models [30, 15, 69, 61, 8, 7, 64] rely on a pretrained vision encoder and, optionally, a text encoder to extract conditioning features. The vision encoders are generally frozen during the training process, serving as visual feature extractors for audio generation. The extracted visual features encompass the global scene context, yet they inherently carry noisy and excessive irrelevant information. Thereby, it impedes the generation of user-intended sound.
@@
We expect hidden video embeddings to contain semantics that are exclusively aligned with text. However, a simple cross-attention mechanism yields a suboptimal result where the model still suffers from generating sounds corresponding to the motion dynamics of non-target instances. For example, when a dog is barking beside a cat, it incorrectly produces a meowing sound, reflecting semantic confusion between co-occurring sound sources. To mitigate this issue, we introduce a novel technique to attach learnable supplementary tokens  [SUP] preceding text embeddings. As shown in Fig. 3 (‘w/o [SUP]’), such artifacts emerge as semantic patch tokens (particularly those associated with motion dynamics) become high-norm outliers in the attention space. However, as in prior approaches on vision transformers [14, 2], adding extra tokens to the sequence of visual embeddings increases the computational cost across all encoder blocks. Furthermore, we should design to learn intent-focused visual representations that emphasize regions and cues relevant to the user’s specified sound source. To this end,  [SUP] are simply prepended to text features as follows:
@@ 3.3 Selective sound generation
Our generator, $\mathcal{G}$, adopts a multimodal diffusion transformer (MM-DiT) architecture [18, 40]:
@@ 3.4 Training
Video-mixing
As input videos usually comprise multiple sound sources, without explicit annotations for each sound source, it is nontrivial to isolate visual features. To address this, we introduce a self-supervised strategy, motivated by audiovisual separation works [17, 41], but reformulated for selective V2A generation. Concretely, two videos are randomly selected, and horizontally concatenated with a random ratio to make the desired {mixed-video, audio, text} pairs. One of the audio-text pairs is randomly chosen to serve as the target. Formally, an input video $V$ in a mini-batch consists of randomly selected two videos $\{V_{\text{tar}},V_{\text{pair}}\}\in\mathbb{R}^{H\times W}$:
@@
Learning conditional feature extraction from mixed sources and multiple conditioned audio generation simultaneously is inherently complex, as both modules depend on each other’s evolving representations [26]. To ensure efficient and stable optimization, the joint training is organized into a two-stage learning scheme, allowing each module to converge toward a consistent representation before mutual conditioning. In the first stage, the video encoder learns to understand text prompts from output features of the teacher model [68, 66]. In the left part of Fig. 2, while the teacher model $\mathcal{F}_{T}$ (i.e., pretrained Synchformer [31]) takes a single source video $V_{\text{tar}}$ to generate a pseudo feature $\mathbf{v}_{\text{tar}}$, the student model $\mathcal{F}_{S}$ produces a text-guided visual feature from the mixed source. Formally, the video feature $\mathbf{v}$ of student encoder $\mathcal{F}_{S}$ is learned to minimize the L2-norm regression loss:
@@
Next, we train a generator $\mathcal{G}$ while keeping the video encoder frozen in the second stage. We start from the MM-DiT in MMAudio [8] as the baseline of the generator. Rather than finetuning the whole parameters, we focus on specific modules that handle video features explicitly, as illustrated in the right part of Fig. 2. Specifically, we finetune two sub-modules only: (1) the initial projection layer of the video feature branch (i.e., $W_{\mathbf{v}}$) and (2) the adaptive LayerNorm (adaLN) module of the audio latent branch of multimodal and single-modal transformer blocks (i.e., $W_{\gamma},W_{\beta}$). The model is trained with conditional flow-matching (CFM) [46, 59]. Given a noise distribution $q(\mathbf{a}_{0})\sim\mathcal{N}(\mathbf{0},I)$, training data distribution $q(\mathbf{a}_{1}=\mathbf{a}_{\text{tar}},\mathbf{c})$ with input condition features $\mathbf{c}=(\hat{\mathbf{v}}_{\text{tar}}=\mathcal{F}_{S}(V,\mathbf{t}_{\text{tar}}),\mathbf{t}_{\text{tar}})$, and timestep $t\in[0,1]$, the CFM objective is formulated as: